OpenStudy (anonymous):

Please help with cofactors. I am confused because it asks twice for the cofactor of 1. 3x3 Matrix, -1 3 6 1 6 1 -3 1 6 I got 35 for the first one, does that sound right?

OpenStudy (anonymous):

I am trying to find the cofactor of 1, 6, and 1

OpenStudy (anonymous):

is this furby town

OpenStudy (phi):

they seem to be asking for the cofactors of the 2nd row 1 6 1 1 is in the (2,1) position (row 2 col 1) and 2+1=3 is odd. so the sign will be negative now select out the submatrix by excluding the first column and the 2nd row you are left with 3 6 (top row, cols 2 and 3) 1 6 (bot row, cols 2 and 3) the det of this submatrix is 3*6-1*6= 18-6= 12 with the minus sign, the co-factor for 1 (in (2,1)) is -12